OpenStudy (anonymous):

-4x-2y=-12 4x+8y=-24 Solve either using Substitution or elimination. Post your work

OpenStudy (anonymous):

First you will want to solve one equation for a single variable like y

OpenStudy (anonymous):

you have two options to solve from -4x-2y=-12 or 4x+8y=-24

OpenStudy (triciaal):

To solve by elimination add the 2 eqns and solve for y with this value of y put in either eqn 1 or eqn 2 to solve for x. to solve by elimination you let the coefficient of one variable be the same in each of the eqns so they will cancel out here the x are already the same so just add them if they were not the same find factors that will make them the same before adding

OpenStudy (anonymous):

Show how you did it. Says that in the post. I don't need an explanation on how to do it. I want to see how others do it.

OpenStudy (triciaal):

to solve by substitution you would find what one variable is in terms of the other then substitute the expression in the other eqn. this will give one eqn with one variable. using this result substitute to find the other. for example with this question using eqn 1 let x = (-12 + 2y )/-4 use this in eqn 2 4x+8y=-24 same as 4(x + 2y)= -6 x + 2y = -6/4 (-12 + 2y)/-4 + 2y = -6/4 simply and solve for y use this value in x = (-12 + 2y)/-4 to get x

OpenStudy (triciaal):

we are here to help you to do it not to do it for you. how will you learn if you are not trying. now pick which method you want to use and follow the directions. if you post your work I will be happy to check it. I am here to work with you not to just provide an answer.

OpenStudy (anonymous):

I have the answers? lol. Im just seeing if others know how to? featherhh your dismissed now

OpenStudy (anonymous):

Bitcccccchhhhhhhhhhhhhhhhh*